if 16 is added to one-third of a number, the result is 3 times the number. What is the number? Does the proper equation start with: 16 + x/3 = 3x?
Answer by rfer(16322)

1/3x+16=3x
-2 2/3x=-16
-8/3x=-16
x=-16*-3/8
x=48/8
x=6